OpenStudy (anonymous):

Derive the equations for (v1x)f=(m1-m2)/(m1+m2)(v1x)f and (v2x)f=2m1/(m1+m2)(v1x)i from the equations m1(v1x)i=m1(v1xf)+m2(v2x)f and 1/2m1(v1x)i^2=1/2m1(v1x)f^2+1/2m2(v2x)f^2

OpenStudy (anonymous):

I got to (v1x)i-m2(v2x)f/m1=(v1x)f and (v2x)f=sqr((m1((vix)i^2-(v1x)f^2)/m2)

OpenStudy (anonymous):

sub the (v2x)f back into (v1x)i-m2(v2x)f/m1=(v1x)f .